OpenStudy (anonymous):

If y varies inversely as x to the 2 power and y = 3 when x = 18, what is y when x = 3?

OpenStudy (anonymous):

y would be in the form: y = K/x^2. Find K. When x = 18 -> y = 3--> K = 3*x^2 = 3*324 = 972. The function is: y = 972/x^2. When x = 3 --> y = 972/9 = 108. Check: when x = 18 -> y = 972/324 = 3.